OpenStudy (anonymous):

A golfer hits a golf ball with a club head angled at 30 degrees. He swings with a velocity of +50.0 m/s. How far will the ball travel horizontally? 220m 160m 50m 130m

OpenStudy (gollywolly):

Determine vertical and horizontal velocities. Horizontal velocity will remain constant so you should just need to figure out how long it will travel up and down then apply the time to horizontal velocity to find distance. \[\cos(\theta)=v_x/v\]\[\sin(\theta)=v_y/v\]

OpenStudy (astrophysics):

This is a projectile motion problem, it's always good to draw a diagram in physics problems  notice we can use the range formula \[\huge x(range) = \frac{ v_0^2\sin(2\theta) }{ g }\]